OpenStudy (anonymous):

Find the initial value problem dy/dx=5x^-14/15 I got this answer wrong on my test need worked showed out step by step do not understand this to much

OpenStudy (paxpolaris):

\[\Large {dy \over dx}=5x^{(-14/15)}\]is this what you mean????

OpenStudy (anonymous):

yes

OpenStudy (paxpolaris):

don't you also need a 2nd piece of information ... y(a)=b or are you just having trouble with the integration?

OpenStudy (anonymous):

yes missing a piece sorry y(-1)=-3

OpenStudy (paxpolaris):

\[\Large y=5\int\limits x^{(-14/15)}dx\]

OpenStudy (anonymous):

yes i get that far but after that i am completly lost

OpenStudy (paxpolaris):

\[\large -{14 \over 15} +1 = {1\over15}\]

OpenStudy (anonymous):

ohh ok i see that part

OpenStudy (paxpolaris):

\[\Large y= 5\cdot {x^{(1/15)}\over 1/15}+C\]

OpenStudy (paxpolaris):

\[\LARGE =75\cdot \sqrt[15]{x}+C\]

OpenStudy (anonymous):

y=75x^1/15+C

OpenStudy (paxpolaris):

yes ... now you just have to plug in x=-1 , y=-3 ... to solve for C

OpenStudy (anonymous):

oh okay i see now i was not understanding the part you just worked out but i understand now

OpenStudy (anonymous):

the final answer is 75x^1/15+72 ???

OpenStudy (paxpolaris):

\[\Huge \checkmark\] y(0)=72

OpenStudy (anonymous):

ok thanks for your help